OpenStudy (anonymous):

Solve for a: 4a/9 = 100/45

OpenStudy (anonymous):

(100/45)/5= 20/9 4a/9=20/9 20/4=5 a=5

OpenStudy (anonymous):

you should put this is the math section! this is english lol. cx